OpenStudy (anonymous):

P(dry followed by dry)=0.85 P(dry followed by wet)=0.15 Probability of dry day is 0.80 Question: whats is the Probability(wet day followed by another wet day)??

OpenStudy (anonymous):

i really need help with this assignment question

OpenStudy (anonymous):

prob of wet day 1-0.8 = 0.2 probability of wet day followed by wet day 0.2*0.2=0.04

OpenStudy (kc_kennylau):

If P(dry)=0.8, then P(dry=>dry) should be 0.8^2=0.64 ...

OpenStudy (anonymous):

im looking at both yours answers and working one says 0.04 and the other 0.64 hmmmm

OpenStudy (kc_kennylau):

No I'm not the answer, I'm just doubting the validity of the question.

OpenStudy (anonymous):

ye the question just doesnt give out much info which isnt any help

OpenStudy (anonymous):

what ya mean by validity?? @kc_kennylau

OpenStudy (kc_kennylau):

The first statement and the third statement contradict each other @divu.mkr

OpenStudy (anonymous):

yep...

OpenStudy (anonymous):

good point :) @kc_kennylau